OpenStudy (anonymous):

how do you simplify [(x-2)/(x+2)]+[(x-1)/(x+1)]/[(x)/(x+1)]-[(2x-3)/(x)]?

OpenStudy (anonymous):

[(x-1)/(x+1)]/[(x)/(x+1)] this one you flip and time= [(x-1)/(x+1)]*[(x+1)/(x)] -(2x^2-5x+30]/(x*(x+2) = [(x-1)*(x+1)]/[x*(x+1] and simplify you have = (x-1)/x now start add and stract [(x-2)/(x+2)]+[(x-1)/x]-[(2x-3)/(x)] denominators is: x(x+2) x*(x-2x)*+[(x+2)*(x-1)]-[(x+2)*(2x-3)] [(x^2-2x)+(x^2+x-2) -(2x^2+x-6)]/(x+(x+2) simlyfly: [x^2-2x+x^2+x-2-2x^2-x+6]/(x*(x+2) (-2x+4)/(x*(x+2) = -2(x-2)/x*(x-2)